OpenStudy (anonymous):

Members of the Ski Club contributed equally to obtain $1800 for a holiday trip. When 6 members found they could not go, their contributions were refunded and each remaining member then had to pay an additional $10 to remain at their goal of $1800. How many went on the trip?

OpenStudy (anonymous):

Let x = original number of members who signed up y = original amount each member paid Equation 1: xy = 1800 Equation 2: (x - 6)(y + 10) = 1800 Expanding Equation 2, xy - 6y + 10x - 60 = 1800 From Equation 1, xy = 1800 and y = 1800/x and therefore, Equation 2 becomes 1800 - 6(1800/x) + 10x - 60 = 1800 -10800/x + 10x - 60 = 0 Multiplying the above by "x", -10800 + 10x^2 - 60x = 0 or x^2 - 6x - 1080 = 0 Using the quadratic equation, x = 36 --- this is the number of original members who signed up Therefore, the number who went to the trip = 36 - 6 = 30 members

OpenStudy (anonymous):

thank you for going in depth of how to do it, it will give me something to go off of.

OpenStudy (anonymous):

Yup No Problem :)